### v2

# Language Models and Logic Programs for Trustworthy Financial Reasoning

Abstract
According to the United States Internal Revenue Service, “the average American spends \$270\\$270 and 13 hours filing their taxes”. Even beyond the U.S., tax filing requires complex reasoning, combining application of overlapping rules with numerical calculations. Because errors can incur costly penalties, any automated system must deliver high accuracy and auditability, making modern large language models (LLMs) poorly suited for this task. We propose an approach that integrates LLMs with a symbolic solver to calculate tax obligations. We evaluate variants of this system on the challenging StAtutory Reasoning Assessment (SARA) dataset, and include a novel method for estimating the cost of deploying such a system based on real-world penalties for tax errors. We further show how combining up-front translation of plain-text rules into formal logic programs, combined with intelligently retrieved exemplars for formal case representations, can dramatically improve performance on this task and reduce costs to well below real-world averages. Our results demonstrate the promise and economic feasibility of neuro-symbolic architectures for increasing equitable access to reliable tax assistance.
Introduction
“GPT is not a certified tax professional, nor am I, so you should always check with your tax advisor.”
— Greg Brockman, CTO of OpenAI
Of life’s two certainties, taxes should be preferred; yet they may well be the more complicated one. Each year, virtually every adult in the world must calculate and pay a fee to some government, in order to reside and earn a living within the state’s guardianship. Even for individuals with relatively simple financial situations, the annual filing process demands meticulous reading and following of dozens of form instructions and the copying of values across schedules, worksheets, and eligibility tests. Completing these tasks without professional assistance can take hours. Alternatively, taxpayers may hire a professional preparer, incurring substantial fees depending on the complexity of their return (Internal Revenue Service 2025).
Accuracy in tax filing is essential. Over-reported income or missed deduction opportunities lead to unnecessary over-payment, while under-reporting may result in penalties, interest, and potential legal consequences. In the United States, the costs of inaccuracies affect lower income communities more significantly, in part because these groups offer the Internal Revenue Service (IRS) high audit success rates (Black et al. 2022; Elzayn et al. 2025). However, these audits deliver a modest return on investment compared to audits of wealthier taxpayers, so there is an opportunity to better align community and institutional interests through improved tax advice to lower income taxpayers (Boning et al. 2024).
Figure 1: A taxpayer confronted with a tax question might choose between an inexpensive AI preparer and a costlier human professional. The decision considers trade-offs between cost, convenience, and confidence in the result.
Figure 2: Methods for solving. Top Left: Plain-text for statutes and a case is fed into a language model, along with the instruction to calculate a person’s tax obligation. Top Right: Statutes and a case are fed into the model as before, but it is instructed to convert these into a logic program which calculates a person’s tax obligation. If the SWI-Prolog engine fails to execute the program, the case is considered unanswered. Bottom: A language model parses a case’s facts into Prolog, conditioned on gold parses of the most relevant cases and of the rules contained in the statutes. The symbolic solver imports the gold parses of the statutes before attempting to execute the generated parse of the case. Note that unlike the approaches above it, this requires gold symbolic representations of both the statutes and a representative selection of correctly-decided cases.
However, the concrete costs for errors present a substantial challenge for modern large language models (LLMs). An AI assistant deployed in this domain must meet higher standards than basic accuracy: it should (1) recognize when it lacks sufficient certainty to offer guidance and (2) generate a transparent and faithful trace of logical steps so that taxpayers and auditors can easily verify the derivation of each answer. In this paper we show that symbolic reasoning tools, integrated with LLMs, offer a promising approach to meeting these standards. Our method provides the language model with access to a symbolic solver, enabling it to translate statutory text and taxpayer information into formal logic programs, which are processed by a trusted execution engine. We evaluate the method on the StAtutory Reasoning Assessment (SARA) dataset, a benchmark of synthetic tax scenarios paired with liability calculations carried out through ground-truth representations of rules and facts in formal logic (Holzenberger et al. 2020).
Our experiments demonstrate two key findings. First, whereas frontier reasoning models outperform non-reasoning models at both directly solving and at parsing case and statute text into the symbolic solver, non-reasoning models consistently outperform their reasoning counterparts when given gold symbolic representations of statutes and of their application to similar cases. Second, we show that by adding additional refusal criteria through a symbolic solver and self-checking, the expected costs of deploying such a system in the real world could be brought down to less than 20% of the average cost for an American to file their taxes. Our results indicate the promise of neuro-symbolic architectures for expanding access to trustworthy and reliable tax expertise.
Background
Logic Programming for Legal Reasoning
Several programming languages have been designed to represent and facilitate logical reasoning. Prolog is a declarative programming language for representing and reasoning over knowledge, with roots in first order logic. A programmer defines rules using Horn clauses (Horn 1951) and facts by declaring which rules apply to entities, thus populating a knowledge base. Subsequently, this knowledge base can be queried by defining a ‘goal’, which launches computation in the form of a backward-chaining search attempting to prove that the goal holds a certain value (Wielemaker et al. 2010). Prolog has been used since the early days of legal AI, where it has formed the backbone of legal expert systems because its declarative syntax keeps knowledge base entries for rules human-readable while powerfully representing the reasoning around which legal questions revolve (Sherman 1989). Efforts in countries like the United Kingdom (Sergot et al. 1986), Canada (Sherman 1987), and the United States (Kant et al. 2025) have leveraged this capacity to encode legal rules in executable formal logic. Related languages have also been used in the legal domain, such as Answer-set Programming (Gelfond and Lifschitz 1988; Morris 2020), Datalog (Ceri et al. 1989; Huang et al. 2021), and Catala (Merigoux et al. 2021a; Merigoux 2023). More broadly, hierarchical templates are a popular tool for evaluation of legal reasoning (Hou et al. 2024). We focus on the SARA dataset (Holzenberger et al. 2020), which encodes statutes and cases into Prolog logic programs to show how a symbolic expert system can perfectly solve a task which large language models struggle to complete.
Figure 3: Number of correct and incorrect solutions produced by each solution method, for large chat- and reasoning-optimized models (served by DeepSeek and OpenAI).
Statutory Tax Reasoning and the SARA Dataset
We focus on the task of statutory reasoning for tax law. Some elements of this task bear similarity to popular mathematical reasoning tasks such as GSM-8k (Cobbe et al. 2021) or MATH-500 (Hendrycks et al. 2021), such as chaining together mathematical operations to solve a real-world problem described in words. However, unlike these math datasets which require application of a small set of universal arithmetic rules which models learn during training, statutory reasoning considers a set of contingent rules contained within documents provided to a model in-context at inference time, in addition to these basic arithmetic principles.
We evaluate our methods on the SARA dataset, which tests the ability of language models to do statutory reasoning about the United States Tax Code (Holzenberger et al. 2020). This dataset is included in the popular aggregate benchmark LegalBench (Guha et al. 2023), and was used in the GPT-4 product launch to highlight the model’s superior reasoning capacity (Blair-Stanek et al. 2024). The SARA dataset consists of 9 sections from the US federal tax code which have been moderately edited to make them self-contained and unambiguous. These manipulations allow the dataset to serve as a self-contained and solvable task for language models that lack live internet access and human-like abilities to process ambiguity (Jurayj et al. 2022; Stengel-Eskin et al. 2024). These statute sections are accompanied by 376 hand-crafted cases to test understanding of these statutes, each containing a question about a person’s tax obligation. Each statute and case has been manually translated into Prolog, which allows them to be trivially solved using the language’s powerful execution engine to resolve queries about cases. This Prolog is defined using neo-Davidsonian event semantics (Davidson 1966), thus categorizing each event as one of 61 possible predicates onto which various arguments are attached. Of the 376 cases and corresponding questions, 276 require binary judgments about whether a statute applies to a given case, and 100 require numerical judgments about how much tax a person owes in a given year. We focus on these 100 tax cases because of their increased difficulty, and because a trivial baseline of always guessing a single answer delivers poor performance at predicting the numerical output.
Zero-Shot Solving from Statutory Text
Direct Calculation
We evaluate our methods against the baseline method of direct solving, mirroring the approach used in OpenAI’s GPT-4 demonstration (Blair-Stanek et al. 2024). This approach treats the tax calculation as a mathematical question answering task with the additional demand that the model must apply entries from a large corpus of rules contained in the statutory text, in addition to the generic arithmetic rules that govern all calculation. For each case in this setting, a model’s context is filled with all sections of the statutes concatenated together, the description of the case’s facts, and the question about a person in that case’s tax liability, all in plain text. It is instructed to calculate the person in question’s tax obligation based on the rules outlined in the statutes.
Chat Model Reasoning Model Size
Qwen2.5-Coder R1-Distill Qwen2.5 32 billion
Llama 3.3 R1-Distill Llama 3.3 70 billion
DeepSeek-V3 DeepSeek-R1 671 billion
GPT-4.1 OpenAI o3 \$8\\$8/m tokens
Table 1: Chat and Reasoning model pairs. Each open-weight model pair is fine-tuned from the same base model. Although the exact dimensions and provenance of OpenAI’s models are unknown, the two models have identical token pricing structures, suggesting that they incur similar costs for OpenAI to serve.
Calculation by Parsing for a Symbolic Solver
To extend the direct solution approach, we augment the language model with a symbolic solver. Here, a model is given the plain text of the statutes as in the direct calculation case. It is instructed to generate a Prolog program which encodes the relevant rules and facts necessary to compute the person in question’s tax obligation. The symbolic solver ingests a set of rules and facts in Prolog, and is invoked to execute a query. The execution of this Prolog program offers a straightforward mechanism for refusal: if the program fails to execute into the proper format or hangs beyond a pre-allocated time limit (10 seconds), the system is considered to have refused to answer.
Model Family Model Method Correct Incorrect Abstentions Break-Even Price
Baseline N/A Always Abstain 0 0 100 \$270 ± 0
N/A Always Predict \$0 5 95 0 \$16227.11 ± 7805.94
Qwen-2.5 Qwen-32b Direct 13 87 0 \$3,051.64 ± 1,828.31
Qwen-32b Parsed 2 17 81 \$490.34 ± 230.75
R1-32b Direct 38 62 0 \$505.25 ± 287.67
R1-32b Parsed 1 2 97 \$278.70 ± 24.33
Llama-3.3 Llama-70b Direct 9 91 0 \$1,065.90 ± 675.07
Llama-70b Parsed 1 43 56 \$252,027.73 ± 414,049.97
R1-70b Direct 43 57 0 \$1,257.03 ± 1,620.47
R1-70b Parsed 2 1 97 \$266.10 ± 6.81
DeepSeek DeepSeek-V3 Direct 22 78 0 \$739.45 ± 474.59
DeepSeek-V3 Parsed 11 43 46 \$2,099.13 ± 1,253.57
DeepSeek-V3 Direct + Direct 16 15 69 \$265.46 ± 63.53
DeepSeek-V3 Direct + Parsed 7 4 89 \$285.53 ± 55.57
DeepSeek-V3 Parsed + Parsed 5 8 87 \$310.47 ± 67.95
DeepSeek-R1 Direct 74 26 0 \$304.29 ± 225.57
DeepSeek-R1 Parsed 38 10 52 \$249.64 ± 84.77
DeepSeek-R1 Direct + Direct 66 12 22 \$94.20 ± 59.76
DeepSeek-R1 Direct + Parsed 34 3 63 \$170.10 ± 21.75
DeepSeek-R1 Parsed + Parsed 17 4 79 \$241.80 ± 29.45
OpenAI GPT-4.1 GPT-4.1 Direct 48 52 0 \$532.84 ± 492.99
GPT-4.1 Parsed 39 31 30 \$228.89 ± 151.69
GPT-4.1 Direct + Direct 42 13 45 \$196.92 ± 88.43
GPT-4.1 Direct + Parsed 27 6 67 \$185.10 ± 21.33
GPT-4.1 Parsed + Parsed 26 5 69 \$186.30 ± 20.84
o3 Direct 56 44 0 \$6,431.84 ± 2,637.94
o3 Parsed 75 15 10 \$47.43 ± 22.16
o3 Direct + Direct 41 17 42 \$3,472.29 ± 1,859.32
o3 Direct + Parsed 52 10 38 \$115.90 ± 24.63
o3 Parsed + Parsed 65 9 26 \$77.51 ± 22.41
OpenAI GPT-5 GPT-5 Direct 76 24 0 \$299.11 ± 288.41
GPT-5 Parsed 53 13 34 \$122.72 ± 29.21
GPT-5 Direct + Direct 73 9 18 \$218.64 ± 270.19
GPT-5 Direct + Parsed 46 6 48 \$138.30 ± 25.53
GPT-5 Parsed + Parsed 31 5 64 \$180.23 ± 23.42
Table 2: Results of different methods without gold statutes. Models in the same family have the same base model (or seem most likely to, in the case of closed-weights models). Note that “break-even price” measures only the costs of failures and abstentions, and does not include inference costs. For each model, the approach that delivered the lowest break-even price is shown in bold. The top two rows show the break-even price of trivial systems, which always defer to an expert or which always tell a person not to pay any taxes. Errors represent a 90% confidence interval. The lowest break-even price method for each model family is in bold.
Experimental Setup
We run experiments across four model families of different sizes, three of which are open-weight models. The bases for these models are: Qwen2.5 32B (Qwen et al. 2025), Llama 3.3 70B (Grattafiori et al. 2024), DeepSeek-V3 671B (DeepSeek-AI et al. 2025b), and OpenAI’s GPT-4.1 (OpenAI 2025a, b), each of which has an instruction-tuned version designed for common chat applications, and a reasoning version optimized to expend additional inference-time compute to solve harder problems. The full list of models is included in table 1. We run auxiliary experiments using GPT-5, but we do not conduct the same chat vs. reasoning comparison because this product appears to be an integrated system containing several models, and therefore is not as analogous to these other comparisons (Zhang et al. 2025). The reasoning model for three of these pairings stem from the DeepSeek R1 project (DeepSeek-AI et al. 2025a), where strong base models were fine-tuned to generate long chains-of-thought that help them solve harder quantitative reasoning problems. Although there is no formal documentation stating that OpenAI’s GPT-4.1 and o3 models are derived from the same base model, the proximity of the two model’s launch dates and their identical per-token pricing schemes suggest that these two models have a similar relationship to the other model pairs we explore.
We consider ‘correct‘ attempts to be those where the output calculated by our system is exactly the same as the actual tax obligation, when rounded to the nearest dollar. All Prolog code is executed using the SWI-Prolog (Wielemaker et al. 2010) implementation of Prolog, and externally halted after 10 seconds of reasoning. We run experiments on the v2 release of the SARA dataset, because its programmatic representations most closely match the natural language surface forms (Holzenberger and Van Durme 2021).
Self-Consistency Tests
We further ask how effectively these methods can serve to improve each other’s selectivity, by using comparisons between different solution methods to expend additional compute to help determine whether an answer should be trusted (Wang et al. 2023; Stengel-Eskin and Van Durme 2023; Jurayj et al. 2025). In these settings, an answer is only accepted if it is reached via two independent reasoning processes: two chains-of-thought and answers (either directly calculated obligations or Prolog programs) are sampled from the same model. When self-checking using the same method (for instance, “Parsed + Parsed”), these answers are conditioned on the same prompt and context. In the parsing-based approaches, this can be considered a more stringent version of the existing refusal system which rejects questions for which the parsed solution does not execute, by additionally deferring to a tax professional where a combination of attempts do not reach a consensus answer. We test the effectiveness of each combination of reasoning processes for each model to show where additional selectivity can further improve performance.
Incorporating Costs of Incorrect Judgments
To file a faulty tax return can incur substantial financial costs. These can take the form of government-imposed penalties for understatement, or simply the cost of paying more in tax than one actually owes. Previous evaluations of SARA typically report exact match scores on the tax cases. Recent work has examined how large the errors are between the obligations calculated by a language model, observing that smaller errors are more frequent than larger ones (Blair-Stanek et al. 2024). We extend this approach by calculating the costs that would be incurred by using these systems to file the taxes for the 100 tax cases in the SARA dataset.
We posit that the deployment of an automated tax advisor system should require some accountability for the organization deploying it. To provide a realistic estimate of the costs of employing the methods we outline above draw from the US Internal Revenue Code (IRC) $\S 6662$, which imposes a penalty of $20\%$ of the amount underpaid for “substantial understatement of income tax”. The cases in SARA are all personal or household tax cases, so a tax filing is considered a “substantial understatement” if they are less than the maximum of $10\%$ of the actual amount owed or \$5,000\\$5,000. We finally impose a penalty of \$270\\$270 for refusing to answer, to model the cost of the personal or professional time required to complete one’s taxes (Internal Revenue Service 2025).
This gives us the following formula: Let $N$ be the number of cases the system ingests, and $\Delta y_{i}=y_{i}-\hat{y}_{i}$ be the difference between the actual obligation $y_{i}$ and predicted obligation $\hat{y}_{i}$, such that positive values of $\Delta y_{i}$ indicate understatement and negative values indicate overstatement.
$\text{cost}=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}-(\Delta y_{i}),&\Delta y_{i}<0,\\[6.0pt] 0.2\,(\Delta y_{i}),&\Delta y_{i}>\max\!\bigl{(}\$5,000,\;0.1\cdot y_{i}\bigr{)},\\[6.0pt] $\$270$,&\text{if refused},\\[6.0pt] $\$0$,&\text{otherwise}.\par\end{cases}$
The first line corresponds to the cost of overstatement (i.e. the amount overstated), the second and fourth lines correspond to the respective fees incurred for substantial and non-substantial understatements, while the third line simulates the average cost for an American to file their taxes according to the IRS (Internal Revenue Service 2025).
This cost also corresponds to the break-even price of the tax assistant, i.e. the minimum price at which they might offer this service without becoming insolvent. This simulates a real-world scenario in which an organization assumes liability for the costs of the errors their system makes for their users, and offers tax filing with deferral to a tax expert at a fixed price point. Here, the break-even price would inform the minimum price at which they might offer this service without becoming insolvent, such that a more accurate system delivers a lower break-even price. LLM inference costs remain under \$1 per-question, and are omitted because of their marginal impact.
We note that the penalty discussed can also be imposed for “negligence or disregard of rules or regulations” while filing taxes. One interpretation might be that using an AI system to complete one’s taxes is inherently negligent, but for the purposes of this work we assume this is not the case.
Exemplars Successes Failures Break-Even Price
Ranked 87 8 \$247.99\\$247.99
Random 71 24 \$5459.25\\$5459.25
Table 3: Intelligently re-ranking few-shot exemplars improves performance. We measure the advantage of intelligent versus random exemplar re-ranking on GPT-4.1.
Few-Shot Parsing using Gold Statutes
We additionally consider what advantage could be given by offering language model parsers access to gold symbolic representations of the statutes, alongside demonstrations of cases applying the rules from these statutes, building on previous work showing how demonstrations can aid in semantic parsing (Shin and Van Durme 2022; Spiegel et al. 2024). Notably, the gold parsed cases all reference the same manually-translated Prolog representation of the statutes. As such, these manual translations formalize certain types of facts in one particular way where several may be viable, given alternative representations of the rules which are plausible but not implemented in practice. Thus, when using already parsed cases as few-shot examples, the execution engine must have access to this particular formalization of the rules as well.
Figure 4: Success and failure rates of method mixtures: The top right corner counts the average number of successes yielded by each method combination, and the bottom left corner counts the average number of failures for models over 100 billion parameters optimized for reasoning (DeepSeek R1 and OpenAI o3) and chat (DeepSeek V3 and GPT-4.1)
To identify the most salient exemplar cases, we apply information retrieval systems which can follow instructions (Weller et al. 2025a), directing a retrieval system to rank cases based on how similar the logical structure of the case’s text is to the case at hand. For each case, we instruct a lightweight reasoning model (OpenAI o4-mini) to rank the other 99 cases, following recent work showing the effectiveness of test-time scaling for reranking documents (Weller et al. 2025b; Yang et al. 2025). As few-shot examples, we provide the 5 most relevant cases and their gold Prolog translations in-context for the language model to condition its parse on. These cases can be analogized to precedents, because our tax calculation agent uses them to understand how terms from the statutes are applied in practice. We note that this use of the term ‘precedent’ is informal and does not refer to the common-law practice of binding precedent. Rather, these retrieved precedent cases are more analogous to ‘persuasive’ precedent, which might help a court understand how terms have been applied in the past but is not itself new law which future courts must apply (Kozel 2014).
To test the value of this intelligent selection process for few-shot examples, we compare the performance of GPT-4.1, which table 4 shows is the strongest model in this setting, against its performance with randomly selected few-shot examples in table 3. Although the model delivers strong performance when conditioned even on random exemplars, its performance is substantially improved through the intelligent exemplar retrieval process.
Model Family Model Method Successes Failures Abstentions Break-Even Price
Previous Best o3 Parsed 75 15 10 \$47.43 ± 22.16
Qwen-2.5 Qwen-32b Few-Shot 42 38 20 \$4,676.49 ± 4,623.77
R1-32b Few-Shot 47 33 20 \$7,783.29 ± 6,244.96
Llama-3.3 Llama-70b Few-Shot 70 27 3 \$1,917.32 ± 1,247.50
R1-70b Few-Shot 29 57 14 \$6,328.48 ± 2,545.70
DeepSeek DeepSeek-V3 Few-Shot 78 18 4 \$468.66 ± 273.47
DeepSeek-V3 Direct + Few-Shot 18 2 80 \$223.43 ± 19.69
DeepSeek-V3 Parsed + Few-Shot 9 1 90 \$250.43 ± 15.31
DeepSeek-V3 Few-Shot + Few-Shot 73 9 18 \$271.45 ± 230.09
DeepSeek-R1 Few-Shot 40 16 44 \$378.73 ± 155.14
DeepSeek-R1 Direct + Few-Shot 32 2 66 \$178.20 ± 21.34
DeepSeek-R1 Parsed + Few-Shot 19 2 79 \$234.65 ± 28.76
DeepSeek-R1 Few-Shot + Few-Shot 20 3 77 \$215.33 ± 20.63
OpenAI GPT-4.1 GPT-4.1 Few-Shot 87 8 5 \$247.99 ± 341.76
GPT-4.1 Direct + Few-Shot 47 0 53 \$143.10 ± 22.49
GPT-4.1 Parsed + Few-Shot 38 1 61 \$164.70 ± 21.98
GPT-4.1 Few-Shot + Few-Shot 81 5 14 \$40.08 ± 15.87
o3 Few-Shot 81 13 6 \$60.26 ± 58.93
o3 Direct + Few-Shot 51 5 44 \$126.41 ± 24.54
o3 Parsed + Few-Shot 68 7 25 \$75.11 ± 22.46
o3 Few-Shot + Few-Shot 74 8 18 \$58.13 ± 20.91
OpenAI GPT-5 GPT-5 Few-Shot 86 9 5 \$15.78 ± 10.34
GPT-5 Direct + Few-Shot 71 5 24 \$64.98 ± 19.23
GPT-5 Parsed + Few-Shot 50 2 48 \$129.60 ± 22.51
GPT-5 Few-Shot + Few-Shot 84 6 10 \$29.28 ± 13.84
Table 4: Results of different methods with access to gold statutes and intelligently retrieved parsing exemplars. Columns have the same meaning as table 2. Note that because these results require the additional work of manually translating all statutes and a set of representative cases, they are not directly comparable to those in table 2. The top row shows the best previous result.
Results
We display the effectiveness of each method which doesn’t use the gold statutes in table 2. To extend these, we show the effectiveness of methods which access gold statutes and exemplars in table 4, although we note that these are not directly comparable to the results in table 2 because they require more up-front human effort.
We observe a substantial divergence between the effectiveness of reasoning- and chat-optimized models on different variants of this task. We visualize this in fig. 3, aggregating performance of the more powerful DeepSeek and OpenAI models optimized for Chat and Reasoning. Figure 4 shows how this disparity is further amplified when an answer must be reached via two independent reasoning paths in order to be accepted; whereas reasoning models can effectively mix methods to check their work, chat models deliver exceptional performance when self-checking few-shot solutions, but weaker performance otherwise.
Discussion
Our results indicate the promise of augmenting large language models with a symbolic solver. In both settings with and without access to gold symbolic representations of the statutes, the most effective method combined the strongest model (i.e. the OpenAI offering) with the Prolog symbolic solver. In addition to this performance advantage, the use of the symbolic solver is desirable for this specific task because it means that taxpayers or auditors may inspect and debug the system’s reasoning process after the fact, with a guarantee that the decision the system reached was achieved by the path which the logic program articulates. Although one might attempt to audit the chains of thought used to solve the problem in the direct solution cases, the causal relationship between these chains and the answer they help produce is less robust than symbolic program execution, and can be deeply misleading to human readers (Paul et al. 2024; Barez et al. 2025; Li et al. 2025; Skaf et al. 2025).
Interestingly, our experiments aggregated in fig. 3 reveal a notable divergence between models optimized for reasoning and for chat applications. Although reasoning models perform better at direct solving and zero-shot parsing into symbolic representations of rules and facts, the chat models exhibit surprising effectiveness at few-shot parsing of case facts. It is possible that reasoning models’ post-training focuses their efforts on emulating explicit reasoning steps, such as those Prolog would itself execute, rather than accurately mapping input texts into symbolic form. Whereas long chains of thought are useful for the complicated arithmetic required for directly calculating a tax obligation or for discerning which portions of the statutes should be included in a zero-shot parse, they degrade performance on the relatively simpler task of imitating exemplary conversions from natural language into formal logic. This could be a manifestation of the “lost in the middle” effect in long-context processing (Liu et al. 2024): when the model is producing its final parse, the case’s plain-text is in between a long block of the statutes’ Prolog and the long chain of thought which the model has just generated, thus distracting the model from the relatively straightforward task of converting from natural language into the neo-Davidsonian Prolog syntax which can comprehensively represent the SARA cases.
Figure 5: Cases with more words are more likely to be miscalculated. We note a moderate correlation between the case length and the failure rate of solution attempts.
In practice, this additionally means that with gold exemplars of case and statute translations, an individual’s tax case could be processed much more quickly than it could be without these manual annotations, because chat models take dramatically less time than reasoning models to produce answers. This could help build more interactive tax assistance tools which allow users to quickly iterate to amend and clarify their relevant tax information. This approach notably reduces the expected real-world cost per successful tax filing, highlighting the critical role that intelligent exemplar selection can play in guiding model outputs toward faithful and structured reasoning. The most effective configuration at minimizing the system’s break-even price is achieved when providing models with intelligently retrieved precedential examples, and deferring to a tax expert unless this system yields the same answer from two independently sampled solutions. The break-even price of \$49.48\\$49.48 for this system would save the average American taxpayer over $80\%$ of the amount spent on tax filing (Internal Revenue Service 2025). This strong performance could be a strong first step towards real-world pilot studies.
Of course, to expect that statutory codes would be manually translated into a logic programming language like Prolog is a constraining assumption to deliver this strong performance. However, we note that a handful of existing organizations have embarked on this task, such as projects to encode the French tax code (Merigoux et al. 2021b) or Canadian policy proposals (Morris 2020) into logic programs. In the United States, private companies build similar logic programs into the backend of consumer tax arrangement software (Yu et al. 2020). As this practice increases in popularity, it will expand opportunities to implement methods which use gold-standard programmatic representations of rules.
We look closely at the cases which are most and least commonly solved correctly by the most powerful models (DeepSeek and OpenAI). We see in fig. 5 that cases which have more words in them are more likely to be mistaken across models, perhaps because they are more likely to contain intricate logical structures which are more difficult to translate. For instance, table 5 shows that both case 91 and case 83 were miscalculated by every approach. These two cases have an unusual predicate structure that does not appear in other cases, involving the person in question hiring overlapping groups of people for different periods of the year. Although either case appears in the other case’s few-shot examples when these are selected by the intelligent reranking system, the single example appears insufficient for any of the models we test to grasp the application of this unusual logical structure. Real-world deployment should consider how additional heuristics for case difficulty can further inform the decision to defer to a tax professional.
Least Frequently Failed Most Frequently Failed
Case Index Errors Case Index Errors
tax_case_89 0 tax_case_26 12
tax_case_11 1 tax_case_47 12
tax_case_76 1 tax_case_10 12
tax_case_45 1 tax_case_91 12
tax_case_56 2 tax_case_83 12
Table 5: Most and least frequently mistaken cases.
Finally, we note in table 2 a dramatic jump in performance with scale at parsing cases and statutes without gold translations. Although smaller reasoning models (those derived from Qwen2.5 32B and Llama 3.3 70B) perform solidly at direct solving, and the smaller chat models perform solidly at few-shot parsing, neither category of smaller model solves more than a handful of cases in the zero-shot parsing setting. In contrast, all larger models solve at least $10\%$ of these cases correctly, while for OpenAI o3 this is the most successful setting. We hope that further increases in model scale will further improve performance in this setting, since it combines the low human cost of direct solving with the easy auditing of systems that rely on the symbolic solver.
Conclusion
We show how the integration of symbolic solvers with frontier language models can enhance the capability of AI tax assistance systems, and highlight their potential to improve equitable access to accurate and affordable tax guidance. Although tradeoffs remain between up-front costs of translating rules into formal logic versus ongoing inference-time computational and error costs, leveraging symbolic reasoning reduces overall expenses while improving auditability in both cases. Future research will explore how further scale or specialized smaller models optimized for faithful translation can improve aspects of this task, and develop methods for efficient translation of statutory rules into formal logic to enable effective few-shot learning. Finally, real-world implementation should be facilitated through interactive user studies using a further optimized system, enabling widespread accessibility to trusted, cost-efficient AI tax assistance.

### v3

@@
-Language Models and Logic Programs for Trustworthy Financial Reasoning
+Language Models and Logic Programs for Trustworthy Tax Reasoning
 Abstract
-According to the United States Internal Revenue Service, “the average American spends \$270\\$270 and 13 hours filing their taxes”.
+According to the United States Internal Revenue Service, “the average American spends $\mathdollar 270$ and 13 hours filing their taxes”.
 Even beyond the U.S., tax filing requires complex reasoning, combining application of overlapping rules with numerical calculations.
@@
 We further show how combining up-front translation of plain-text rules into formal logic programs, combined with intelligently retrieved exemplars for formal case representations, can dramatically improve performance on this task and reduce costs to well below real-world averages.
-Our results demonstrate the promise and economic feasibility of neuro-symbolic architectures for increasing equitable access to reliable tax assistance.
+Our results demonstrate the effectiveness of applying semantic parsing methods to statutory reasoning, and show promising economic feasibility of neuro-symbolic architectures for increasing access to reliable tax assistance.
+Code is available at https://github.com/wjurayj/legal_logic_programs
 Introduction
@@
 Over-reported income or missed deduction opportunities lead to unnecessary over-payment, while under-reporting may result in penalties, interest, and potential legal consequences.
-In the United States, the costs of inaccuracies affect lower income communities more significantly, in part because these groups offer the Internal Revenue Service (IRS) high audit success rates (Black et al.
+In the United States, the costs of inaccuracies affect lower income communities more significantly, in part because these groups offer the Internal Revenue Service (IRS) high audit success rates (Dean 2022; Black et al.
 2022; Elzayn et al.
@@
 2020).
-Our experiments demonstrate two key findings.
+Our experiments are the first to demonstrate strong performance by applying semantic parsing methods to statutory reasoning.
+This includes two key findings.
 First, whereas frontier reasoning models outperform non-reasoning models at both directly solving and at parsing case and statute text into the symbolic solver, non-reasoning models consistently outperform their reasoning counterparts when given gold symbolic representations of statutes and of their application to similar cases.
@@
 A programmer defines rules using Horn clauses (Horn 1951) and facts by declaring which rules apply to entities, thus populating a knowledge base.
-Subsequently, this knowledge base can be queried by defining a ‘goal’, which launches computation in the form of a backward-chaining search attempting to prove that the goal holds a certain value (Wielemaker et al.
-2010).
+Subsequently, this knowledge base can be queried by defining a ‘goal’, which launches computation in the form of a backward-chaining search attempting to prove that the goal holds a certain value by iteratively unifying sub-goals (Wielemaker et al.
+2010; Weir et al.
+2024).
 Prolog has been used since the early days of legal AI, where it has formed the backbone of legal expert systems because its declarative syntax keeps knowledge base entries for rules human-readable while powerfully representing the reasoning around which legal questions revolve (Sherman 1989).
@@
 Each statute and case has been manually translated into Prolog, which allows them to be trivially solved using the language’s powerful execution engine to resolve queries about cases.
-This Prolog is defined using neo-Davidsonian event semantics (Davidson 1966), thus categorizing each event as one of 61 possible predicates onto which various arguments are attached.
-Of the 376 cases and corresponding questions, 276 require binary judgments about whether a statute applies to a given case, and 100 require numerical judgments about how much tax a person owes in a given year.
-We focus on these 100 tax cases because of their increased difficulty, and because a trivial baseline of always guessing a single answer delivers poor performance at predicting the numerical output.
-Zero-Shot Solving from Statutory Text
+This Prolog is defined using neo-Davidsonian event semantics (Davidson 1966), categorizing each event as one of 61 possible predicates onto which various arguments are attached.
+276 of these cases require binary responses about whether a section applies, and the remaining 100 require calculating a tax obligation; we focus on the 100 cases because of their increased difficulty and because a trivial baseline of always guessing a single answer delivers poor performance at predicting the numerical output.
+Methodology
 Direct Calculation
@@
 DeepSeek-V3 DeepSeek-R1 671 billion
-GPT-4.1 OpenAI o3 \$8\\$8/m tokens
+GPT-4.1 OpenAI o3 $\mathdollar 8$/m tokens
 Table 1: Chat and Reasoning model pairs.
@@
 Although the exact dimensions and provenance of OpenAI’s models are unknown, the two models have identical token pricing structures, suggesting that they incur similar costs for OpenAI to serve.
-Calculation by Parsing for a Symbolic Solver
+Zero-Shot Parsing for a Symbolic Solver
 To extend the direct solution approach, we augment the language model with a symbolic solver.
 Here, a model is given the plain text of the statutes as in the direct calculation case.
-It is instructed to generate a Prolog program which encodes the relevant rules and facts necessary to compute the person in question’s tax obligation.
-The symbolic solver ingests a set of rules and facts in Prolog, and is invoked to execute a query.
+It is instructed to generate a Prolog program which encodes the relevant rules and facts necessary to compute the person in question’s tax obligation, reframing the task from reasoning to autoformalization (Zhang et al.
+2025a).
+The symbolic solver ingests a set of rules and facts in Prolog, and is invoked to execute a query, offloading to the solver the compositional reasoning that language models struggle to conduct faithfully (Khandelwal and Pavlick 2025).
 The execution of this Prolog program offers a straightforward mechanism for refusal: if the program fails to execute into the proper format or hangs beyond a pre-allocated time limit (10 seconds), the system is considered to have refused to answer.
@@
 The top two rows show the break-even price of trivial systems, which always defer to an expert or which always tell a person not to pay any taxes.
-Errors represent a 90% confidence interval.
+Errors represent a symmetricized 90% confidence interval.
 The lowest break-even price method for each model family is in bold.
+Few-Shot Parsing using Gold Statutes
+We additionally consider what advantage could be given by offering language model parsers access to gold symbolic representations of the statutes, alongside demonstrations of cases applying the rules from these statutes, building on previous work showing how demonstrations can aid in semantic parsing (Shin and Van Durme 2022; Spiegel et al.
+2024).
+Notably, the gold parsed cases all reference the same manually-translated Prolog representation of the statutes.
+As such, these manual translations formalize certain types of facts in one particular way where several may be viable, given alternative representations of the rules which are plausible but not implemented in practice.
+Thus, when using already parsed cases as few-shot examples, the execution engine must have access to this particular formalization of the rules as well.
+This reduces the per-case reasoning task to tax-relevant event extraction (Holzenberger and Van Durme 2023; Gantt et al.
+2024; Walden et al.
+2025)
+To identify the most salient exemplar cases, we apply information retrieval systems which can follow instructions (Weller et al.
+2025a), directing a retrieval system to rank cases based on how similar the logical structure of the case’s text is to the case at hand.
+For each case, we instruct a lightweight reasoning model (OpenAI o4-mini) to rank the other 99 cases, following recent work showing the effectiveness of test-time scaling for reranking documents (Weller et al.
+2025b; Yang et al.
+2025).
+As few-shot examples, we provide the 5 most relevant cases and their gold Prolog translations in-context for the language model to condition its parse on.
+These cases can be analogized to precedents, because our tax calculation agent uses them to understand how terms from the statutes are applied in practice.
+We note that this use of the term ‘precedent’ is informal and does not refer to the common-law practice of binding precedent.
+Rather, these retrieved precedent cases are more analogous to ‘persuasive’ precedent, which might help a court understand how terms have been applied in the past but is not itself new law which future courts must apply (Kozel 2014; Savelka 2020; Šavelka and Ashley 2022).
 Experimental Setup
@@
 reasoning comparison because this product appears to be an integrated system containing several models, and therefore is not as analogous to these other comparisons (Zhang et al.
-2025).
+2025b).
 The reasoning model for three of these pairings stem from the DeepSeek R1 project (DeepSeek-AI et al.
@@
 We further ask how effectively these methods can serve to improve each other’s selectivity, by using comparisons between different solution methods to expend additional compute to help determine whether an answer should be trusted (Wang et al.
-2023; Stengel-Eskin and Van Durme 2023; Jurayj et al.
-2025).
+2023; Stengel-Eskin and Van Durme 2023).
 In these settings, an answer is only accepted if it is reached via two independent reasoning processes: two chains-of-thought and answers (either directly calculated obligations or Prolog programs) are sampled from the same model.
@@
 We test the effectiveness of each combination of reasoning processes for each model to show where additional selectivity can further improve performance.
+Figure 4: Success and failure rates of method mixtures: The top right corner counts the average number of successes yielded by each method combination, and the bottom left corner counts the average number of failures for models over 100 billion parameters optimized for reasoning (DeepSeek R1 and OpenAI o3) and chat (DeepSeek V3 and GPT-4.1)
 Incorporating Costs of Incorrect Judgments
@@
 To provide a realistic estimate of the costs of employing the methods we outline above draw from the US Internal Revenue Code (IRC) $\S 6662$, which imposes a penalty of $20\%$ of the amount underpaid for “substantial understatement of income tax”.
-The cases in SARA are all personal or household tax cases, so a tax filing is considered a “substantial understatement” if they are less than the maximum of $10\%$ of the actual amount owed or \$5,000\\$5,000.
-We finally impose a penalty of \$270\\$270 for refusing to answer, to model the cost of the personal or professional time required to complete one’s taxes (Internal Revenue Service 2025).
+The cases in SARA are all personal or household tax cases, so a tax filing is considered a “substantial understatement” if they understate the true obligation by more than the maximum of $10\%$ of the actual amount owed or $\mathdollar 5,000$.
+We finally impose a penalty of $\mathdollar 270$ for refusing to answer, to model the cost of the personal or professional time required to complete one’s taxes (Internal Revenue Service 2025).
 This gives us the following formula: Let $N$ be the number of cases the system ingests, and $\Delta y_{i}=y_{i}-\hat{y}_{i}$ be the difference between the actual obligation $y_{i}$ and predicted obligation $\hat{y}_{i}$, such that positive values of $\Delta y_{i}$ indicate understatement and negative values indicate overstatement.
-$\text{cost}=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}-(\Delta y_{i}),&\Delta y_{i}<0,\\[6.0pt] 0.2\,(\Delta y_{i}),&\Delta y_{i}>\max\!\bigl{(}\$5,000,\;0.1\cdot y_{i}\bigr{)},\\[6.0pt] $\$270$,&\text{if refused},\\[6.0pt] $\$0$,&\text{otherwise}.\par\end{cases}$
+$\text{cost}=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}-(\Delta y_{i}),&\Delta y_{i}<0,\\[6.0pt] 0.2\,(\Delta y_{i}),&\Delta y_{i}>\max\!\bigl(\mathdollar 5,000,\;0.1\cdot y_{i}\bigr),\\[6.0pt] $\textdollar 270$,&\text{if refused},\\[6.0pt] $\textdollar 0$,&\text{otherwise}.\par\end{cases}$
 The first line corresponds to the cost of overstatement (i.e.
@@
 One interpretation might be that using an AI system to complete one’s taxes is inherently negligent, but for the purposes of this work we assume this is not the case.
-Exemplars Successes Failures Break-Even Price
-Ranked 87 8 \$247.99\\$247.99
-Random 71 24 \$5459.25\\$5459.25
-Table 3: Intelligently re-ranking few-shot exemplars improves performance. We measure the advantage of intelligent versus random exemplar re-ranking on GPT-4.1.
-Few-Shot Parsing using Gold Statutes
-We additionally consider what advantage could be given by offering language model parsers access to gold symbolic representations of the statutes, alongside demonstrations of cases applying the rules from these statutes, building on previous work showing how demonstrations can aid in semantic parsing (Shin and Van Durme 2022; Spiegel et al.
-2024).
-Notably, the gold parsed cases all reference the same manually-translated Prolog representation of the statutes.
-As such, these manual translations formalize certain types of facts in one particular way where several may be viable, given alternative representations of the rules which are plausible but not implemented in practice.
-Thus, when using already parsed cases as few-shot examples, the execution engine must have access to this particular formalization of the rules as well.
-Figure 4: Success and failure rates of method mixtures: The top right corner counts the average number of successes yielded by each method combination, and the bottom left corner counts the average number of failures for models over 100 billion parameters optimized for reasoning (DeepSeek R1 and OpenAI o3) and chat (DeepSeek V3 and GPT-4.1)
-To identify the most salient exemplar cases, we apply information retrieval systems which can follow instructions (Weller et al.
-2025a), directing a retrieval system to rank cases based on how similar the logical structure of the case’s text is to the case at hand.
-For each case, we instruct a lightweight reasoning model (OpenAI o4-mini) to rank the other 99 cases, following recent work showing the effectiveness of test-time scaling for reranking documents (Weller et al.
-2025b; Yang et al.
-2025).
-As few-shot examples, we provide the 5 most relevant cases and their gold Prolog translations in-context for the language model to condition its parse on.
-These cases can be analogized to precedents, because our tax calculation agent uses them to understand how terms from the statutes are applied in practice.
-We note that this use of the term ‘precedent’ is informal and does not refer to the common-law practice of binding precedent.
-Rather, these retrieved precedent cases are more analogous to ‘persuasive’ precedent, which might help a court understand how terms have been applied in the past but is not itself new law which future courts must apply (Kozel 2014).
-To test the value of this intelligent selection process for few-shot examples, we compare the performance of GPT-4.1, which table 4 shows is the strongest model in this setting, against its performance with randomly selected few-shot examples in table 3.
-Although the model delivers strong performance when conditioned even on random exemplars, its performance is substantially improved through the intelligent exemplar retrieval process.
 Model Family Model Method Successes Failures Abstentions Break-Even Price
@@
 Whereas long chains of thought are useful for the complicated arithmetic required for directly calculating a tax obligation or for discerning which portions of the statutes should be included in a zero-shot parse, they degrade performance on the relatively simpler task of imitating exemplary conversions from natural language into formal logic.
-This could be a manifestation of the “lost in the middle” effect in long-context processing (Liu et al.
-2024): when the model is producing its final parse, the case’s plain-text is in between a long block of the statutes’ Prolog and the long chain of thought which the model has just generated, thus distracting the model from the relatively straightforward task of converting from natural language into the neo-Davidsonian Prolog syntax which can comprehensively represent the SARA cases.
-Figure 5: Cases with more words are more likely to be miscalculated. We note a moderate correlation between the case length and the failure rate of solution attempts.
+This aligns with concurrent work highlighting cases in which long chains of thought can distract models from tasks where shorter chains would deliver stronger performance (Gema et al.
+2025).
+Figure 5: Standalone parsing success rate improves with model size.
+While smaller reasoning models fail to solve nearly every problem in this setting, larger reasoning models show improvement with model size.
 In practice, this additionally means that with gold exemplars of case and statute translations, an individual’s tax case could be processed much more quickly than it could be without these manual annotations, because chat models take dramatically less time than reasoning models to produce answers.
@@
 The most effective configuration at minimizing the system’s break-even price is achieved when providing models with intelligently retrieved precedential examples, and deferring to a tax expert unless this system yields the same answer from two independently sampled solutions.
-The break-even price of \$49.48\\$49.48 for this system would save the average American taxpayer over $80\%$ of the amount spent on tax filing (Internal Revenue Service 2025).
-This strong performance could be a strong first step towards real-world pilot studies.
+The break-even price of $\mathdollar 15.78$ for this system would save the average American taxpayer over $90\%$ of the amount spent on tax filing (Internal Revenue Service 2025).
+This strong performance could be a strong first step towards real-world pilot studies, or on developing open-source model alternatives to GPT-5 that can adaptively allocate computational resources based on problem difficulty (Jurayj et al.
+2025; Wang et al.
+2025, 2026).
 Of course, to expect that statutory codes would be manually translated into a logic programming language like Prolog is a constraining assumption to deliver this strong performance.
@@
 As this practice increases in popularity, it will expand opportunities to implement methods which use gold-standard programmatic representations of rules.
-We look closely at the cases which are most and least commonly solved correctly by the most powerful models (DeepSeek and OpenAI).
-We see in fig.
-5 that cases which have more words in them are more likely to be mistaken across models, perhaps because they are more likely to contain intricate logical structures which are more difficult to translate.
-For instance, table 5 shows that both case 91 and case 83 were miscalculated by every approach.
-These two cases have an unusual predicate structure that does not appear in other cases, involving the person in question hiring overlapping groups of people for different periods of the year.
-Although either case appears in the other case’s few-shot examples when these are selected by the intelligent reranking system, the single example appears insufficient for any of the models we test to grasp the application of this unusual logical structure.
-Real-world deployment should consider how additional heuristics for case difficulty can further inform the decision to defer to a tax professional.
-Least Frequently Failed Most Frequently Failed
-Case Index Errors Case Index Errors
-tax_case_89 0 tax_case_26 12
-tax_case_11 1 tax_case_47 12
-tax_case_76 1 tax_case_10 12
-tax_case_45 1 tax_case_91 12
-tax_case_56 2 tax_case_83 12
-Table 5: Most and least frequently mistaken cases.
-Finally, we note in table 2 a dramatic jump in performance with scale at parsing cases and statutes without gold translations.
+Furthermore, we believe the relative strength of o3 at zero-shot parsing (which includes parsing a subset of the statutory rules) suggests promise in semi-automated offline encoding of these rules; as models improve further at this type of complex semantic parsing, entire legal systems could be distilled into programmatic logic.
+We note in fig.
+5 a dramatic jump in performance with scale at parsing cases and statutes without gold translations.
 Although smaller reasoning models (those derived from Qwen2.5 32B and Llama 3.3 70B) perform solidly at direct solving, and the smaller chat models perform solidly at few-shot parsing, neither category of smaller model solves more than a handful of cases in the zero-shot parsing setting.
@@
 We hope that further increases in model scale will further improve performance in this setting, since it combines the low human cost of direct solving with the easy auditing of systems that rely on the symbolic solver.
+Additionally, this capability could also align with improved performance on formalizing statutes alone, allowing tighter feedback for policy makers on the effects of proposed amendments on government revenue (Blair-Stanek et al.
+2026).
 Conclusion
-We show how the integration of symbolic solvers with frontier language models can enhance the capability of AI tax assistance systems, and highlight their potential to improve equitable access to accurate and affordable tax guidance.
+We frame statutory reasoning as a semantic parsing task, instructing language models to convert statutory and case text into executable code.
+We show how the integration of symbolic solvers with frontier language models can enhance the capability of AI tax assistance systems, and highlight their potential to improve access to accurate and affordable tax guidance.
 Although tradeoffs remain between up-front costs of translating rules into formal logic versus ongoing inference-time computational and error costs, leveraging symbolic reasoning reduces overall expenses while improving auditability in both cases.
-Future research will explore how further scale or specialized smaller models optimized for faithful translation can improve aspects of this task, and develop methods for efficient translation of statutory rules into formal logic to enable effective few-shot learning.
-Finally, real-world implementation should be facilitated through interactive user studies using a further optimized system, enabling widespread accessibility to trusted, cost-efficient AI tax assistance.
+Future research will explore how further scale or specialized smaller models optimized for faithful translation can improve aspects of this task, and develop methods for efficient translation of statutory rules into formal logic to enable effective few-shot learning and help inform policy design.
+Ethical Statement
+Although we believe our results are promising, this paper should not be understood as a recommendation for individuals filing their taxes.
+Instead, we hope that policymakers and software developers will extend these insights to reduce the individual effort required for tax filing, building systems which can allow a single human-in-the-loop to complete filings more efficiently.
+Such steps would require careful consideration of liability issues, impacts on vulnerable populations, and non-financial costs associated with tax reviews such as stress or time.
+Furthermore, the reliance on closed-source models may place private companies at a choke point, so further advances in the open-source domain are essential.
+Acknowledgments
+This work was funded in part by the U.S.
+National Science Foundation under grant No.
+2204926, and by the Defense Advanced Research Project Agency (DARPA) CODORD program.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or DARPA.